Grade 12Mechanics

A rigid ball of mass m strikes a rigid wall at 60 degree and gets reflected without any loss of speed. The value of impulse imparted by the wall on the ball will be ?
  • mv
  • 2mv
  • mv/2
  • mv/3

Abina George
9 Years agoGrade 12

3 Answers

Vikas TU
9 Years ago
Dear Student,
After doing components,
v1= v X root3/2 i^ - v X root3/2 j^
v2= - v X root3/2 i^ - v X root3/2 j^

Impulse on wall = mv1 – mv2
=m root3 x v.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
SUBHAM GHOSH
7 Years ago
Impulse = F × ∆t 
 Now , F = (ma) 
            F = { m(v2 - v1)/ ∆t } 
Impulse = m(v2 - v1)
 
v1 = v cos 60° (i^) + v sin 60° (- j^)
v2 = v cos 60° (- i^) + v sin 60° (- j^)
 
Impulse = m × ( v2- v1 ) 
               = m × 2v cos60° (- i^)
               = m × 2v × ( 1/2) (- i^)
               = mv (- i^)
Kushagra Madhukar
5 Years ago
Dear student,
Please find the attached solution to your problem.
 
As we know, angle of incidence =  angle of reflection.
Now, v1 = v cos 60° i + v sin 60° j    [ where, v is the speed of ball]
Hence, velocity after reflection,
v2 = – v cos 60° i + v sin 60° j
Now, by impulse momentum theorem we have,
Impulse = change in momentum
            = mv2 – mv1
            = m( – 2v cos 60o i ) = – mv i
Hence, magnitude of impulse = mv
 
Hope it helps.
Thanks and regards,
Kushagra